ammonia has a Kb of 1.8x10^-5. find [H3O+], [OH-], pH, and pOH for a 0.390 M ammonia solution
Answer #1

NH3 + (aq) + H2O ( l) <---> NH4+(aq) + OH- (aq)

initially [NH3] =0.39 , [NH4+] =[OH-] = 0

at equilibrium [NH3] = 0.39-X , [NH4+]=[OH-] = X

Kb = [NH4+][OH-]/[NH3]

1.8 x 10^ -5 = (X)(X)/ ( 0.39-X)

X = 0.00264 =[OH-]

[H3O+] = 10^ -14 / [OH-] = 10^ -14 / ( 0.00264) = 3.79 x 10^ -12 M

pH = -log [H3O+] = -log ( 3.79x10^ -12) = 11.42

pOH = 14-11.42 = 2.58